OpenStudy (anonymous):

Find an equation of the tangent line 1/2 {x log(x^2), {-1, 0}}

hartnn (hartnn):

do you know how to find slope of tangent line ?

OpenStudy (anonymous):

I need to take the derivative using the product rule but having trouble

OpenStudy (anonymous):

I get 1/2(ln(x^2)) + 1/2(x)1/x^2(2x) Don't know if that is right,

hartnn (hartnn):

you'll need chain rule to differentiate log (x^2) yes! thats correct :)

OpenStudy (anonymous):

ok cool. The other person that tried to help me said it was wrong. lol

OpenStudy (anonymous):

So from that I would need to somehow get the slope and then use y - y1 = m(x - x1)

hartnn (hartnn):

simplify it .

OpenStudy (anonymous):

well it would be 1/2(ln(x^2) + (x)1/x^2(2x)

OpenStudy (anonymous):

and then perhaps 1/2ln(x^2) + 1/x^2(2x^2)

hartnn (hartnn):

for 2nd term. 1/2 and 2 gets cancelled 2 'x's in numerator gets cancelled with x^2 in denominator. can you see it ?? so, 2nd term is just =1

OpenStudy (anonymous):

so does that change to 1/2 ln(x^2) + 2 ?

OpenStudy (anonymous):

And then how do I get the slope out of that?

OpenStudy (anonymous):

wait, would I plug in the -1? into x?

OpenStudy (anonymous):

If I did that I would get 1/2ln(-1^2) + 2

OpenStudy (anonymous):

and then 1/2 ln(1)+2

OpenStudy (anonymous):

hmm, do you have any suggestions?

OpenStudy (anonymous):

So can someone tell me if this is right? (y-0) = (1/2ln(1)+2)(x-(-1).

OpenStudy (anonymous):

yeah and I missed ) at the end

OpenStudy (anonymous):

hey Tech-Hunter do you understand the problem?

hartnn (hartnn):

sorry, i was disconnected. 1/2(x)1/x^2(2x) simplifies to +1 and not +2. so,slope =1/2 ln(1)+1 is correct. so, equation will be y = (1/2ln(1)+1)(x+1).

hartnn (hartnn):

1/2(x)1/x^2(2x) simplifies to +1 did u understand this ?

hartnn (hartnn):

\[\dfrac{1}{2}x\dfrac{1}{x^2}2x=\dfrac{2}{2}\dfrac{x.x}{x^2}=1\]